OpenStudy (anonymous):

solve the polynomial equation by factoring 3x^5+6x^4-72x^3=0

hero (hero):

\[3x^3(x^2+2x-24) = 0\]

hero (hero):

3x^3 = 0, so x = 0 is one of the solutions

hero (hero):

x^2 +2x - 24 = 0 x^2+ 6x -4x -24 = 0 x(x+ 6)-4(x+6) = 0 (x+6)(x-4)=0

hero (hero):

So the factors are 3x^3(x+6)(x-4) = 0

hero (hero):

And x = -6,0,4

hero (hero):

I hope any of that made sense

OpenStudy (anonymous):

it did. thank you